the measures of the angles in a triangle are in the ratio 2:3:4. describe how to use algebra to find the measures of the angles. then find the measure of each angle.
Answer by checkley77(12569)

First we add the ratios 2+3+4=9
Now we find the multiplier for each ratio.
90/9=10 is the multiplier.
2*10=20
3*10=30
4810=40
Proof:
20+30+40=90
90=90